### v1

# Topological excitations in a spin-orbit coupled spin-1 Bose-Einstein condensate under sinusoidally varying magnetic fields

Abstract
We present a theoretical study of a spin-orbit coupled spin-1 Bose-Einstein condensate under the influence of sinusoidally varying magnetic fields. In the ground state of the ferromagnetic spin-1 condensate, we investigate topological excitations in the system arising due to the combined effect of Rashba spin-orbit coupling and an in-plane sinusoidally varying magnetic field. In this work, we offer a comparative study for various choices of magnetic fields in the $x\rm{-}y$ plane. For a fixed field strength, the spin-orbit coupled system sustains a rich variety of exotic vortex structures ranging from vortex-antivortex lattices to vortex clusters as we increase the coupling strength.
1 Introduction
The experimental realization of spin-orbit coupling (SOC) in ultracold atomic physics [1, 2, 3, 4, 5, 6, 7, 8, 9] stands as a transformative milestone, opening new avenues for exploration at the quantum frontier. This groundbreaking achievement propels the field into uncharted territories where the delicate interplay between particle spins and their orbital motion can be precisely controlled and manipulated [10, 11, 12]. Typically, the order parameter of the spin-orbit (SO) coupled trapped Bose-Einstein condensates (BECs) [3] exhibit a very rich variety of exotic quantum states such as vortex lattice [13, 14], half-quantum vortex [3, 15], vortex neclace [16], skyrmion [17, 18, 19], domain wall [2], solitons [20, 21, 22], knots [23, 24], supersolid stripe phase [25, 26, 27], spin and field squeezing [28, 29] etc. Moreover in recent times, research has been extended in new directions such as SO-induced supersolidity [30, 27], spin-tensor momentum coupling [31], spin-orbital angular momentum coupling [32], etc. The introduction of such couplings has led to many interesting quantum states with novel topological textures [33, 34]. Recently, the splitting of vortex cores [35] and the first-order phase transitions [36] have been experimentally observed. These topological defects in ultracold condensates are vital for drawing analogies with those of condensed matter systems.
In this light, spinor condensates [37, 38, 11, 39] emerge as exceptional platforms within ultracold atomic physics, serving as a prolific area of research for the investigation of solid-state-like properties [40]. These remarkable quantum systems, exhibiting a multitude of spin degrees of freedom, provide an intriguing bridge between the microscopic world of ultracold atoms and the condensed matter physics of solids. Spinor condensates offer a controlled environment to explore phenomena such as magnetic ordering [41], topological excitations [21] and defects [42], and quantum phase transitions [43, 44], shedding light on the intricate connection between quantum gases and condensed matter systems. This convergence of ultracold atomic physics and solid-state physics not only enriches our understanding of fundamental physics but also holds the promise of unveiling new quantum phenomena with potential applications in quantum information processing and beyond [45].
On the other hand, recent literature shows that in-plane magnetic fields [46, 47, 48, 49, 50, 51, 52, 53, 54, 55] can be significantly used in investigating the equisite many-body physics of ultracold atoms. Relevant studies demonstrate that in-plane gradient magnetic fields can be used to achieve SOC and spin Hall states in BECs [56]. These magnetic fields play a vital role in creating synthetic non-Abelian gauge fields [57, 58]. Varying magnetic fields can be accredited for the observation of a wide assortment of topological defects, such as magnetic monopoles [50, 59], quantum knots [23], skyrmions [18], etc. The ground state of spinor BECs in the gradient field has been investigated to exhibit the central Mermin-Ho vortex with symmetrical vortex lattices [60]. The combined effect of SOC and gradient magnetic field on such systems resulted in skyrmion chains [18, 61] in the ground state order parameter. To the best of our knowledge, the effect of a sinusoidally varying magnetic field on SO-coupled BECs has not been given due attention. In this work, we study the combined effect of a sinusoidally varying magnetic field and SOC on the ground state of a spin-1 BEC. Here, we consider an SO-coupled spin-1 BEC subject to various combinations of sinusoidally varying magnetic fields in th $x\rm{-}y$ plane. We offer a comparative study between the various choices of sinusoidally varying magnetic fields, thereby highlighting the different topological structures so obtained for each case.
This paper is organised as follows: In section 2, we introduce our model system and discuss its mean-field ground state under different situations for the different combinations of sinusoidally varying in-plane magnetic field. In section 3, we discuss the results obtained by numerically solving the Gross-Pitaevskii (GP) equations and investigate the topological structures (vortices) generated in the system as a consequence of the interplay between SOC, sinusoidally varying magnetic field with the spin-1 order parameter. In section 4, we summarize our findings and provide an outreach for future scope.
2 Model
In the mean-field, we consider a quasi-2D Rashba SO-coupled ferromagnetic spin-1 BEC of $N$ identical atoms in an harmonic trap in a sinusoidally varying magnetic field. The effective Hamiltonian [62] in the Gross-Pitaevskii form can be given as $H=H_{0}+H_{int}$,
$H_{0}=\int d\mathbf{r}\Psi^{\dagger}\left[-\frac{\hbar^{2}\nabla^{2}}{2M}+V(\mathbf{r})+\nu_{\rm{soc}}+g_{F}\mu_{B}\mathbf{B(\mathbf{r}).f}\right]\Psi$ (1)
[…]
The Rashba SOC interaction [1, 3] is given as $\nu_{SOC}=\gamma(f_{x}p_{y}-f_{y}p_{x})$ where $\gamma$ is the SOC strength and $(p_{x},p_{y})$ represents the momentum in the quasi-2D space. Note that this SOC is SU(2) type, therefore the states $\Psi_{1}$ and $\Psi_{-1}$ cannot be coupled directly. In real experiments, the SU(2) SOC Hamiltonian can be achieved by Raman coupling three laser beams of different polarizations and frequencies. The external magnetic field is given as $B(\mathbf{r})=(B_{x}\hat{x}-B_{y}\hat{y})$ in the 2D plane. Here, $B_{x},B_{y}$ are sinusoidally varying magnetic fields along the $x$ and $y$ directions. We absorb the negative sign in the direction and denote $B(\mathbf{r})=(B_{x},B_{y})$. The Lande’s g-factor $g_{F}=-\frac{1}{2}$ and $\mu_{B}$ is the Bohr’s magneton.
The ground state and dynamics of the system can be described by the following non-dimensionalised coupled GP equations,
$\rm{i}\frac{\partial{\psi_{1}}}{\partial{t}}$ $=\left[-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\left(|\psi_{1}|^{2}+|\psi_{0}|^{2}-|\psi_{\rm{-}1}|^{2}\right)\right]\psi_{1}+c_{2}\psi_{\rm{-}1}^{*}\psi_{0}^{2}$
$+(B_{x}+\rm{i}B_{y})\psi_{0}-\rm{i}{\gamma}\left(\frac{\partial}{\partial y}+\rm{i}\frac{\partial}{\partial x}\right)\psi_{0}$ (3)
$\rm{i}\frac{\partial{\psi_{0}}}{\partial{t}}$ $=\left[-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\left(|\psi_{1}|^{2}+|\psi_{\rm{-}1}|^{2}\right)\right]\psi_{0}+2c_{2}\psi_{1}\psi_{0}^{*}\psi_{\rm{-}1}$
$+(B_{x}-\rm{i}B_{y})\psi_{1}+(B_{x}+\rm{i}B_{y})\psi_{\rm{-}1}-\rm{i}{\gamma}\left(\frac{\partial}{\partial y}-\rm{i}\frac{\partial}{\partial x}\right)\psi_{1}-\rm{i}{\gamma}\left(\frac{\partial}{\partial y}+\rm{i}\frac{\partial}{\partial x}\right)\psi_{\rm{-}1}$ (4)
$\rm{i}\frac{\partial{\psi_{\rm{-}1}}}{\partial{t}}$ $=\left[-\frac{1}{2}\nabla^{2}+V+c_{0}|\psi|^{2}+c_{2}\left(-|\psi_{1}|^{2}+|\psi_{0}|^{2}+|\psi_{\rm{-}1}|^{2}\right)\right]\psi_{\rm{-}1}+c_{2}\psi_{1}^{*}\psi_{0}^{2}$
$+(B_{x}-\rm{i}B_{y})\psi_{0}-\rm{i}{\gamma}\left(\frac{\partial}{\partial y}-\rm{i}\frac{\partial}{\partial x}\right)\psi_{0}$ (5)
where $\psi_{m}=N^{-1/2}l\Psi_{m}\quad(m=\pm 1,0)$ denotes the dimensionless $m$th component wave function and $|\psi|^{2}=\sum_{m}|\psi_{m}|^{2}$ is the total particle density confined in dimensionless external potential $V=(x^{2}+y^{2})/2$. The contact interaction parameters in the dimensionless form are given as $c_{0}=2N\sqrt{2\pi\lambda}(a_{0}+2a_{2})/3l$ and $c_{2}=2N\sqrt{2\pi\lambda}(a_{2}-a_{0})/3l$, respectively. $B,{\gamma}$ are the dimensionless magnetic field and SOC strengths, respectively. In our numerical calculations, the lengths, time, energy(interaction and SOC) and magnetic field are measured in the units of $l,1/\omega_{\perp},\hbar\omega_{\perp}$ and $(g_{F}\mu_{B}l)/\hbar\omega_{\perp}$, respectively.
The vorticity [63] of the $m-$th component can be measured as
$\Omega_{m}=\nabla\times J_{m}$ (6)
with the probability current density,
$J_{m}=\frac{\rm{i}\hbar}{2M}(\psi_{m}\nabla\psi^{*}_{m}-\psi^{*}_{m}\nabla\psi_{m})$ (7)
The topological excitation of the sytem arising due to the formation of vortices is characterised by the spin texture [64]. For a spin-1 BEC, the spin texture can be written as [65, 66],
$S_{\alpha}=\sum_{m,n=0,\pm 1}\psi^{*}_{m}(f_{\alpha})_{m,n}\psi_{n}/|\psi|^{2}\quad(\alpha=x,y,z)$ (8)
The topological charge density which characterises the topological structure of the system is given as,
$q(\mathbf{r})=\frac{1}{4\pi}\mathbf{s}\cdot\left(\frac{\partial\mathbf{s}}{\partial x}\times\frac{\partial\mathbf{s}}{\partial y}\right)$ (9)
and the total topological charge $Q$ is defined as
[…]
Here $\mathbf{s}=S/|S|$ and the topological charge $|Q|$ is unchanged under any exchange of $S_{x},S_{y},S_{z}$.
3 Results and Discussions
We investigate the ground state structures of the spin-orbit coupled ferromagnetic spin-1 BEC under the influence of external in-plane magnetic field. In this section, we first discuss the effect of a polarized magnetic field in the $x$ or $y$ direction. The ground state density profile and the spin textures are shown for cases both with and without SOC. Thereafter we discuss cases for sinusoidally varying in-plane magnetic fields. For each combination of external magnetic field $B(\mathbf{r})=(B_{x},B_{y})$, we present the ground state structures with and without SOC.
We offer a detailed numerical study to analyze the effects of system parameters on the ground state structures. The luxury of the present system lies in the large number of tunable parameters, which includes the spin-independent and spin-exchange contact interactions, SOC strength, magnetic field strength. To focus our attention on the effects of SOC and in-plane magnetic field, we have considered $10^{4}$ number of $\rm{{}^{87}{Rb}}$ atoms confined in a quasi-2D harmonic trap $V(\mathbf{r})$ with transverse trapping frequency, $\omega_{\perp}=2\pi\times 20$ Hz with trap aspect ratio, $\lambda=\omega_{z}/\omega_{\perp}=20$. The $s\rm{-}$wave scattering lengths for channels of total spin 0 and 2 are $a_{0}=101.8a_{B}$ and $a_{2}=100.4a_{B}$ respectively, where $a_{B}$ is the Bohr radius. With these parameters [67], using the split-step Crank-Nicolson scheme [68, 69, 70], we numerically solve the governing GP equations (2-2) for imaginary time propagation to generate the ground state solutions for diverse choices of magnetic fields for numerous SO coupling strengths. All of our simulations run from a spatial extent of $-20l$ to $20l$ in both $x$ and $y$ directions with $2001\times 2001$ grid points. The employed spatial discretization refers to $\Delta x=\Delta y=0.02l$, with time step of numerical integration $\delta t=(2\times 10^{-4})/\omega_{\perp}$.
3.1 Polarized gradient magnetic field
[…]
Figure 1: The ground state densities and momentum distributions for the spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC. The first and second columns show the density ($a1-a3$) and momentum ($b1-b3$) distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components for the $x-$ polarised magnetic field $B=B_{0}x$. The third and fourth columns represent the density ($c1-c3$) and momentum ($d1-d3$) distributions for the $y-$ polarised magnetic field $B=B_{0}y$. The magnetic field strength $B_{0}=0.25$ for both cases. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
Now, with SOC in play, our investigation centers to unravel the competition between the SOC and external magnetic field. Depending on the choices of the external magnetic field applied, the ground state structures for the SO-coupled spin-1 BEC are different. We show the ground state structures for the SOC strength $\gamma=0.5$ for the $x-$polarised magnetic field in figure 2. The ground state densities $n_{\pm 1},n_{0}$ exhibit modulations along the $x=0$ line as shown in Fig. 2 ($a1-a3$). Our numerical simulations show that the formation of vortices in the three components is pertinent to the interplay of SOC and the external magnetic field. The SOC interacts with the magnetic field in such a way that the effective magnetic field is modified. In addition to this, as the SOC couples the intrinsic spin to the particles’ motion, it imparts some extra angular momentum to the system which results in the formation of vortices in the system. Although there are no saddle points in the magnetic field, these vortices appear at the $x=0$
Figure 2: The ground state structures for the spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the the $x-$ polarised magnetic field $B=B_{0}x$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
Figure 3: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 2. The unit length is the oscillator length $l$.
line under the influence of SOC. The vorticity profile associated to each component is shown in Fig. 2 ($b1-b3$). The vortices are slightly deformed along the $y-$direction as the component densities are modulated along the $x=0$ line. The k-space density profiles (as shown in Fig. 2 (c1-c3)) depict that apart from the central peaks, some extra momenta predominantly along the $k_{x}-$direction is observed. These extra peaks at $k_{x}\neq 0$ corresponds to the density modulations along the $x-$ direction. This is due to the fact that SOC influences the distribution of particles in the k-space, similar to the impact of SOC as seen in the distribution of electronic states as a function of momentum in solid-state physics [71].
The combined effect of the in-plane direction dependent gradient field and the SOC yield intricate spin textures in the spin-1 BEC as shown in Fig. 3 ($b$) with corresponding topological charge distribution $q$ shown in Fig. 3($a$). We can identify the presence of a skyrmion at the center of the vortex core, from the spin texture. The local amplification of the skyrmion spin texture is shown in Fig. 3 ($c$).
For the SO-coupled spin-1 BEC in $y-$polarised gradient magnetic field, we can analogously understand the density modulations along the $y=0$ line for all the three components. However, these density modulations converge into the formation of anti-vortices in the system, and the corresponding spin textures have anti-skyrmions along the $x-$axis. It is interesting to note that the state which had fragments in the absence of SOC, now has more vortices or anti-vortices as we observe for the case of $\gamma=0.5$. As we increase the SOC strength $\gamma$, the number of vortices increases. However, for the polarised magnetic field, they appear in the same alignment, with the vortex spacing being reduced.
3.2 Cross-polarised sinusoidally varying in-plane magnetic fields
Thus far, we have discussed how impactful the polarised gradient field can be on the ground state structure of a SO-coupled spin-1 BEC. To further exploit this direction-dependent interplay of external magnetic field and SOC, we now choose sinusoidally varying in-plane magnetic fields in both $x$ and $y$ directions. In general, to make the magnetic field divergenceless, we also apply a magnetic field $B_{z}=-B_{0}zf(x,y)$ in the $z$-direction, where $f(x,y)$ are sinusoidal functions in the $x-y$ plane. However, as we are in the Quasi-2D geometry, for tight binding along the $z-$direction, $B_{z}$ will not affect the spinor condensate. In this section, we center our focus only on the in-plane sinusoidal magnetic field. Experimentally, a polarised sinusoidal field can be realised using a similar technique discussed in [18], shown for the magnetic field in the vertical direction. Campbell et al has also demonstrated a similar in-plane sinusoidal field in their work [72]. We discuss below, various combinations of $B(\mathbf{r})=(B_{x},B_{y})$ in different cases.
3.2.1 Case-\@slowromancapi@ Sine-Sine magnetic field $B(\mathbf{r})=B_{0}(\rm{sin}x,\rm{sin}y)$
For the present case, we choose the magnetic fields as sine functions in both directions, i.e., $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{sin}y$ with equal magnetic strength $B_{0}=0.25$ in both directions. We first present the ground state structures for the no SOC case in Fig. 4. The density profiles $n_{1},n_{0}$ and $n_{-1}$ of the condensates $m_{F}=1,m_{F}=0$ and $m_{F}=-1$, respectively are shown in Fig. 4 ($a1-a3$). The sine-sine type of magnetic field results in the density modulations in $m_{F}=\pm 1$ components with the appearance of topological exciations in the form of a triad of alternating vortex-anti-vortex along both $x-$ and $y-$axes. This is due to the fact that the applied magnetic field from both directions, generate certain saddlepoint structures such that the nature of the triads in $m_{F}=1$ and $m_{F}=-1$ components are complementary to each other. The opposite nature of the triads is due to the fact that the magnetic field appears as conjugate for the $m_{F}=1$ and $m_{F}=-1$ components in the GP equations (2-2). Furthermore, combining with the spin-mixing term, the magnetic field yields vortex brights in the $m_{F}=0$ component at
Figure 4: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the sine-sine magnetic field $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{sin}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
the loci (saddle points of $B(\mathbf{r})$) of vortices in the $m_{F}=\pm 1$ components so as to conserve the overall angular momentum of the spin-1 BEC in the absence of SOC. Thus, the $m_{F}=0$ component has no vortices, instead a high density peak is observed at the center of the condensate. The vorticity profiles associated to the condensate densities are depicted in Fig. 4 ($b1-b3$). These vortices or anti-vortices are coreless as there is no phase singularity in the total density of the system. The k-space density profiles (see Fig. 4 ($c1-c3$)) support observed density modulations in the system. For the $m_{F}=1$ and $m_{F}=-1$ components, apart from the central peak, we observe finite elongations in both the $k_{x}$ and $k_{y}$ directions which correspond to the density modulations with wavelength inversely $\propto k_{x},k_{y}$.
Since, in the absence of SOC, the overall angular momentum of the spin-1 system is conserved, the corresponding topological charge of the system is $|Q|=0$.
In the presence of SOC, the ground state structures emerge as a result of the interplay
Figure 5: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-sine magnetic field $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{sin}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
among the contributions from the magnetic field, spin-orbit coupling, and the spinor order parameter. For SOC strength, $\gamma=0.5$, we show the density profiles of the $m_{F}=1,m_{F}=0$ and $m_{F}=-1$ components in Fig. 5 ($a1-a3$). The system experiences an induced anisotropy as a result of rotational symmetry breaking due to SOC. This induces distinctive ground state structures, giving rise to the formation of unique configurations such as vortices, skyrmions, and unconventional spin textures. The vorticity profiles (see Fig. 5 ($b1-b3$)) depicts the existence of vortices, anti-vortices in the spin-1 condensate. Due to the induced anisotropy, the vortices or anti-vortices are not only deformed in shape, but also biased in opposite axes for the $m_{F}=1$ and $m_{F}=-1$ components, respectively. The $m_{F}=0$ component experiences density modulations in both directions as a consequence of the spin-exchange interaction, SOC and external magnetic field.
The k-space density structures (as shown in Fig. 5 ($c1-c3$)) are also anisotropically
Figure 6: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 5. The unit length is the oscillator length $l$.
Figure 7: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the cosine-cosine magnetic field $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
elongated in the preferred direction of high densities, thereby justifying the density modulations, arising due to the combined outcome of SOC and the sine-sine magnetic field. Analogous to the previously discussed situations, here also we can associate a density modulation with wavelength $\lambda_{i}$ for additional momentum peak at $k_{i}\neq 0,\quad i=(x,y)$.
We show for the above discussed case, the topological charge distribution in Fig. 6 ($a$) and associated spin texture in Fig. 6 ($b$). The overall spin texture has both skyrmions and anti-skyrmions along the two orthogonal axes. Fig. 6 ($c-d$) are the local amplification of the spin textures showing a skyrmion and anti-skyrmion, respectively.
3.2.2 Case-\@slowromancapii@ Cosine-Cosine magnetic field $B(\mathbf{r})=B_{0}(\mathrm{cos}x,\mathrm{cos}y)$
Here, we choose the magnetic fields as cosine functions in both directions, i.e., $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with equal magnetic strength $B_{0}=0.25$ in both directions. Analogously, we first attempt to understand the ground state structures due to this magnetic field independent of the SO-coupling. The component-wise density profiles are shown in Fig. 7 ($a1-a3$). Much likely the sine-sine magnetic field case, we observe for this case also, due to the interaction of magnetic fields in both directions, a certain stable saddlepoint structure is formed, owing to which a combination of equal vortices and anti-vortices are being formed in the $m_{F}=1$ and $m_{F}=-1$ components such that the overall conservation of angular momentum is obeyed in the spin-1 system. The spin-exchange interaction along with the effect of magnetic field, yields vortex brights in the $m_{F}=0$ component. These local high densities in the $m_{F}=0$ component compensate for the vortices or anti-vortices in the adjacent two components. The vorticity profiles associated to the component densities are presented in Fig. 7 ($b1-b3$), which clearly explains the angular momentum conservation in the system. The topological excitations arise in the system in the form of exotic vortex-anti-vortex configuration at the saddle points of the external magnetic field. The density modulations due to such exotic structures can be explained by the k-space density profiles (see Fig. 7 ($c1-c3$)). For this case also, apart from the central peak, there exists extra momenta both in $k_{x},k_{y}$ conforming the density modulations with wavlength inversely $\propto$ non-zero momentum peaks in both directions. The topological charge is zero for the non SOC case to respect the angular momentum conservation.
Now, we discuss further the effect of the cosine-cosine magnetic field on the SO-coupled spin-1 BEC. For a weak SOC with strength $\gamma=0.5$, the rotational symmetry of the system breaks giving rise to certain topological excitations in the system. The density profiles of the SO-coupled spin-1 BEC are shown in Fig. 8 ($a1-a3$). Due to the symmetry-breaking, an anisotropy is induced in the system which renders a very unique density distribution to three components of the spin-1 BEC. Trivially, due to the spin-mixing interaction, the $m_{F}=1$ and $m_{F}=-1$ components complement each other, with the presence of vortices in one at the loci of high density in the other component. Moreover, the vortices in one are compensated by the anti-vortices in the other. In the $m_{F}=0$ component, criss-cross vortex and vortex-brights are observed at the loci of local high densities at the $m_{F}=\pm 1$ components, as a consequence of the density modulations due to the combined effect of SOC, external magnetic field and spin-mixing dynamics in the system. However, these vortices in the $m_{F}=0$ component have opposite vorticity in the stable state configuration. The vorticity profiles as shown in Fig. 8 ($b1-b3$) gives a clear picture about the topological excitations in the system. In Fig. 8 ($c1-c3$), we present the k-space densities of the three components of the spin-1 BEC. Here also, apart from the central peak, extra momenta along the $k_{x},k_{y}$ directions are observed. These corresponds to the density modulations. Interestingly, for the $m_{F}=0$ component, the k-space density is modulated obliquely as the local high densities are in that direction.
We present the associated topological charge distribution in Fig. 9 ($a$) and the spin texture is shown in Fig. 9 ($b$). For the diagonal vortices in the $m_{F}=0$ component, as the cores are symmetric, the skyrmion and anti-skyrmion structures are circular. However, for elongated
Figure 8: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the cosine-cosine magnetic field $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
Figure 9: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 8. The unit length is the oscillator length $l$.
Figure 10: The ground state structures for spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms without SOC in the sine-cosine magnetic field $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
vortices or anti-vortices, their corresponding topological charge distributions are also elliptical. We show in Fig. 9 ($c$), the local amplification of the spin texture for the region of the diagonal skyrmion and anti-skyrmion.
It is paramount to note that in the absence of SOC, the ground state structures for the $sine-sine$ case and the $cosine-cosine$ case can be analogously explained by taking into account the fact that the topological excitations are observed for certain saddle points in the magnetic fields. In this light, it becomes interesting to study the ground state structures for cross-fields, sine-cosine and/or cosine-sine fields.
3.2.3 Case-\@slowromancapiii@ Sine-Cosine magnetic field $B(\mathbf{r})=B_{0}(\rm{sin}x,\rm{cos}y)$
In this case, we study the ground state structures for the spin-1 BEC with the external magnetic field chosen as $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with equal magnetic strength $B_{0}=0.25$. For
Figure 11: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-cosine magnetic field $B_{x}=B_{0}\mathrm{sin}x$ and $B_{y}=B_{0}\mathrm{cos}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
this combination of magnetic fields, in the absence of SOC, the ground state structures exhibit density modulations due to the resultant interaction of the sine-cosine magnetic fields in the $x,y$ directions, respectively. We present the ground state density profiles of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components in Fig. 10 ($a1-a3$). In the absence of SOC, although the overall angular momentum is conserved, signature of topological excitations in the form of vortices and anti-vortices are observed in the $m_{F}=1$ and $m_{F}=-1$ components. These are arranged in pairs of alternating vorticity along the $y-$direction and as alternating vortex-anti-vortex triads in the $x-$direction. Due to different magnetic fields in different directions, the local high densities in the $m_{F}=0$ component (which arrise as an outcome of the spin-exchange interaction of $m_{F}=0$ component with the $m_{F}=\pm 1$ components), are anisotropic in nature. The vorticity profile (see Fig. 10 ($b1-b3$)) confirms the complementary response of the $m_{F}=1$ and $m_{F}=-1$ components to the external magnetic field. The momentum distribution is shown in Fig. 10 ($c1-c3$). For this case also, along with the central peak, extra momentum both in $k_{x},k_{y}$ directions is observed. However, as their magnitudes are not same in orthogonal directions, the density modulations have different wavelengths along different axes.
In the presence of spin-orbit coupling, the system experiences symmetry breaking, due to which an anisotropy is introduced in the spin-1 system under the influence of magnetic field. For this case, the three components undergo density modulations in both directions, however differently for different components. The density as well as vortex distributions are highly anisotropic in nature due to the combined effect of SOC-induced anisotropy and different sinusoidal fileds in different directions. We present the component-wise density profiles in Fig. 11 ($a1-a3$). For the $m_{F}=1$ component, there is a local high density surrounded by vortices in the system, while for the $m_{F}=-1$ component, there is a vortex surrounded by adjacent local densities. The $m_{F}=0$ component has a unique configuration for which a local high density is observed, which is surrounded by anti-vortices along the $x-$direction and an adjacent vortex along the $y-$direction. The corresponding vorticity profiles are given in Fig. 11 ($b1-b3$). For this case also, the coreless vortices have deformed elliptical shape as a consequence of induced anisotropy in the system. The momentum distributions as shown in Fig. 11 ($c1-c3$) show the effect of anisotropy in the density modulations in the spin-1 system. The k-space densities are elongated in opposite directions for the $m_{F}=1$ and $m_{F}=-1$ components, respectively, i.e, for the $m_{F}=1$ case, the central peak is slightly elongated along the $x-$axis while for the $m_{F}=-1$ component the central density peak is slightly elongated in the $y-$direction. This is to respect the density modulations with different wavelengths in the anisotropic system. In addition to this, the extra momenta in the $k_{x},k_{y}$ directions are not equivalent. For the $m_{F}=0$ component, the momentum distribution is symmetric however not the same about the $k_{x},k_{y}$ axis.
The corresponding topological charge distribution is shown in Fig. 12 ($a$). The presence of skyrmions and anti-skyrmions is confirmed by the spin texture of the system shown in Fig. 12 ($b$). Because of the combined effect of anisotropy and different magnetic fields along different directions, the spin texture has both skyrmions and anti-skyrmions. Fig. 12 ($c$) shows the local amplification of the spin texture showcasing the central vortex-anti-vortex triad.
Figure 12: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 11. The unit length is the oscillator length $l$.
3.2.4 Case-\@slowromancapiv@ Cosine-Sine magnetic field $B(\mathbf{r})=B_{0}(\rm{cos}x,\rm{sin}y)$
For the present case, we propose an alternate choice of cross-magnetic field in the $x-y$ plane, i.e., $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{sin}x$. Here, we attempt to highlight the striking similarities and subtle differences of this present case with the previous case (discussed in section 3.2.3). It is trivial to understand that for this case also, the ground state structures are direction-dependent
Figure 13: The ground state structures for spin-orbit coupled spin-1 BEC of $10^{4}$ $\rm{{}^{87}{Rb}}$ atoms with SOC strength $\gamma=0.5$ in the sine-cosine magnetic field $B_{x}=B_{0}\mathrm{cos}x$ and $B_{y}=B_{0}\mathrm{sin}y$ with strength $B_{0}=0.25$. The first column ($a1-a3$) shows the density distributions, second column ($b1-b3$) shows the vorticity profiles and the third column ($c1-c3$) shows the momentum distributions of the $m_{F}=1$, $m_{F}=0$ and $m_{F}=-1$ components. The momentum space densities have been normalised by dividing the maximum value. The unit length is the oscillator length $l$.
as a result of choosing different magnetic fields in different directions. Importantly, as the magnetic fields along the $x,y-$directions are swapped, their effect on the spin-1 ground state will therefore, be also swapped. In the absence of SOC, the density profiles, associated vorticities and momentum space densities of the three components are interchanged along the $x-y$ axes.
However, for the SO-coupled spin-1 BEC in the cosine-sine magnetic field, the ground state structures are not simply the transpose of that of the sine-cosine case. The ground state structure which is an outcome of the intricate interplay between the external magnetic field, SOC and the spinor order parameter is being presented in Fig. 14 for the SOC strength $\gamma=0.5$. The
Figure 14: ($a$) shows the topological charge distribution, ($b$) shows the spin texture of the system and ($c$) shows the local amplification of the spin texture for the SO-coupled spin-1 BEC shown in Fig. 13. The unit length is the oscillator length $l$.
SOC results in rotational symmetry-breaking of the system rendering the system an induced anisotropy. The density profiles (see Fig. 13 ($a1-a3$)) vouch for the experienced anisotropy in the system. The $m_{F}=1$ component has a unique configuration of vortices along with adjacent local high densities in the system. The $m_{F}=0$ component has a linear chain of vortices in the $y-$direction with an additional anti-vortex in the $x-$direction. The $m_{F}=-1$ component has a local high density surrounded by vortex structures on adjacent sides. The vorticity profile as shown in Fig. 13 ($b1-b3$) depicts the nature of this vortices or anti-vortices in the system. The coreless vortices have a slightly distorted shape due to the symmetry breaking. The momentum distributions are presented in Fig. 13 ($c1-c3$). Due to the induced anisotropy in the system, the k-space densities have visible elongations along a particular direction. In addition, there exists certain extra momenta in the $k_{x},k_{y}$ direction which is not symmetric.
The topological charge distribution is shown in Fig. 14 ($a$). The skyrmions and anti-skyrmions are observed in different axes. The spin texture is shown in Fig. 14 ($b$). Aside from the central triad formed with alternating vortex-anti-vortex configuration, a few other spin textures can be observed owing to the coreless vortices in the system. Fig. 14 ($c$) shows the local amplification of the spin texture showing the central triad in the system.
4 Conclusion
In conclusion, the interplay between spin-orbit coupling and an external in-plane magnetic field with sinusoidal variation in the $x-y$ direction in a spin-1 BEC introduces an additional layer of complexity to an already intricate system. The sinusoidal variation of the magnetic field in the $x-y$ plane provides a unique spatial modulation that further influences the ground state of the condensate. The topological excitations in the form of vortex formation occurs at the saddle points of the in-plane magnetic field. Along with the spin-exchange interaction and spin-mixing dynamics, the magnetic field exhibits conjugate vortices and anti-vortices in the $m_{F}=\pm 1$ components and vortex brights in the $m_{F}=0$ component. These density modulations correspond to a wavelength $\lambda$ which is inversely proportional to the non-zero momentum values in the k-space densities.
In this work, we have attempted to showcase numerous cases with various combinations of sinusoidal fields in the $x-y$ direction. We have discussed the ground state structures for all of these choices both with and without the SOC, highlighting the subtle differences in each case. The SOC induces an anisotropy in the system which breaks the rotational symmetry of the system. As a consequence of which, the overall angular momentum is not conserved in the system. Furthermore, the combination of SOC and the spatially varying magnetic field gives rise to spatially modulated spin textures, including the formation of unique skyrmions and anti-skyrmions. These topological defects represent fascinating quantum states within the condensate, showcasing the intricate interplay between the periodicity induced by the sinusoidal magnetic field and the coupling of spin and motion.
The sinusoidal variation introduces periodicity into the system, creating an opportunity to engineer and manipulate the condensate’s properties. This perocidity of magnetic field can be exploited to see the effect of various combinations of in-plane magentic fields. Beyond this, the dynamics of a SO-coupled spin-1 BEC in a polarised sinusoidal magnetic field can be explored. Moreover, it will be interesting to look into the skyrmion dynamics of the SO-coupled system. The ability to tune the sinusoidal variation of the magnetic field in a controlled environment not only provides a deeper understanding of the quantum phases emerging from the interplay but also opens up possibilities for designing novel quantum devices and applications. As investigations into this intricate interplay continue, we anticipate the discovery of new spatially modulated quantum states, offering exciting prospects for both fundamental research and technological advancements.
Acknowledgements
We thank Hari Sadhan Ghosh for many insightful discussions. We acknowledge the National Supercomputing Mission (NSM) for providing computing resources of PARAM Shakti at IIT Kharagpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India.
Funding information
A.S. gratefully acknowledges the support from the Prime Minister’s Research Fellowship (PMRF), India. S.H. acknowledges the MHRD, Govt. of India for the research fellowship. S.D. acknowledges support from AFOSR FA9550-23-1-0034.